# QUITE: A Query Rewrite System Beyond Rules with LLM Agents

### v1

@@ 4. LLM Agent-Based Query Rewriter
In this section, we present our LLM agent–based query rewriter and describe how to construct an efficient, robust query‐rewriting workflow using LLM agents.
@@ 4.1. MDP-based Reasoning Agent
Query rewrite is fundamentally a sequential decision problem (Finance and Gardarin, 1991b; Zhou et al., 2021), each transformation step changes the query structure and limits future optimization options. Reasoning LLM agents (e.g., DeepSeek-R1 (Guo et al., 2025)), capable of chain-of-thought planning and multi-step logical inference, appear well-suited to this task. However, our experiments show they still produce 13.0% nonequivalent outputs and yield minimal execution improvements (Section 7.2). We observe that query rewrite mirrors a Markov Decision Process, where states are query forms, actions are rewrite steps, and rewards are cost reductions (Puterman, 1990; Liu et al., 2023; Wang, 2025). By framing rewrite in this way, we provide the reasoning LLM agent with a clear, stepwise optimization goal that guides each inference step. Therefore, we design an MDP-based reasoning agent to produce high-performing rewrites.
@@ 4.2. FSM-based Query Rewrite
Single-agent approaches often produce suboptimal or invalid rewrites because a single agent is responsible for managing the entire rewrite pipeline (e.g., propose rewrites, verify validity, and assess efficiency), which overloads the model’s context capacity and ultimately increases the risk of hallucination. Inspired by multi-agent frameworks that assign specialized agents to distinct tasks (Hong et al., 2023), we break down query rewrite into discrete stages, each handled by a dedicated LLM agent. To ensure the multi-agent system’s stability and correctness (Han et al., 2024), we implement a Finite State Machine (FSM) to orchestrate stable stage transitions.
@@ 4.2.1. Query Rewrite Process Decomposition
In practice, we decompose the rewrite process into four stages: Reasoning, Verification, Decision and Termination, as illustrated in Figure 7. The Reasoning stage generates candidate SQL rewrites. The Verification stage validates syntax and semantic equivalence, correcting any errors identified. The Decision stage determines whether to continue or terminate the rewrite process.
@@ 29       $t\leftarrow t+1$;
Rewrite Agent. After the Reasoning agent generates a chain, the Rewrite agent extracts proposals from each node, groups them into four categories matching our structured knowledge base, evaluates each proposal’s expected reward, and selects the top SQL candidate. Since this candidate may introduce new structural complexities, it cannot be used directly for verification. Therefore, the Rewrite agent performs an SQL Enhancement Process—using the prompt template below—to identify further rewrite opportunities.
@@ 3. Apply other efficiency-preserving transformations
  • Equivalence correction. Once syntax‐correct, the agent submits the query pair $\langle Q_{\text{orig}},Q_{\text{new}}\rangle$ to verify the equivalence and correct nonequivalent queries.
@@
(2) Evaluation and Feedback Loop. Based on the generated report, the Decision agent gives a binary judge to decide whether the rewrite process should terminate or not. If True, the Decision agent terminates the FSM process; If False, it retrieves higher-confidence rewrite proposals from the knowledge base, then stores the current query, report, and retrieved knowledge in an agent shared buffer for the next FSM iteration. We deliberately avoid introducing knowledge base retrieval in the initial state, because reasoning models have input content limitations—excessive guidance information can reduce the generation freedom, potentially compromising the quality of reasoning proposal chains (Guo et al., 2025). Finally, the FSM enters the Termination stage and outputs the final rewritten query along with its comprehensive optimization report.
@@ 5.1. Structured Knowledge Base
Preparing the knowledge base involves the following steps: (i) collecting SQL rewrite knowledge from various resources, (ii) filtering untrustworthy content, (iii) enhancing the knowledge, and (iv) retrieving relevant knowledge for specific tasks. The Knowledge Base $\mathcal{K}$ is defined as follows:
@@
Step 3: Enhancing the knowledge. An LLM first summarizes questions and user’s responses from website Q&A units. These summaries are then augmented with corroborating information from official documents. To do this, we embed user response summaries using Sentence Transformers (Reimers and Gurevych, 2019). These embeddings are then used to query a pre-indexed knowledge base of extracted points, retrieving the top three matches via cosine similarity. Finally, an LLM evaluates these matches to confirm they effectively explain the rewrite strategy underlying the Q&A unit. Suitable matches are incorporated, enriching Q&A units with concise explanations that aid future retrieval.
Step 4: Retrieving relevant knowledge. We classify these units into four categories: join optimization, constant folding, predicate simplification and other, which capture the essential aspects of query rewriting and provides a foundational structure for enabling precise retrieval of information. Since knowledge bases are often context-specific, we use the BM25 algorithm (Robertson et al., 2009) to retrieve the most relevant documents by computing the document relevance.
@@ 5.2. Hybrid SQL Corrector
Equivalence Error Correction. Verifying the equivalence between any pair of SQL queries is known to be an NP-hard challenge (Aho et al., 1979), and whether this problem admits a polynomial-time solution remains an open question. While numerous verification methods (Chu et al., 2017a; Wang et al., 2022, 2024; Ding et al., 2023) rooted in algebraic or symbolic reasoning have been proposed to prove equivalence, they typically operate on constrained query sets (e.g., specific operators, advanced SQL constructs). Recent studies have highlighted the potential of LLMs in deducing query equivalence with high confidence (Singh and Bedathur, 2024; Zhao et al., 2023). However, relying solely on LLMs carries inherent risks. Therefore, we propose a hybrid approach that integrates traditional verification tools with LLM capabilities to achieve broader and more reliable equivalence judgments in two stages.
@@ 5.3. Agent Memory Buffer
No Hash Join /*+ NoHashJoin(table table[ table...]) */ Prevents the use of hash joins for specified tables.
@@
NO_MATERIALIZE /*+ NO_MATERIALIZE(table) */ Inlines the CTE to avoid materialization overhead.
@@ 6. Query Hint Recommender
Motivation. Query rewrite applies high-level transformations to give the DBMS optimizer a better starting point for plan generation. While this improved starting point is beneficial, the optimizer, which dictates the finer-grained execution operations, can still generate suboptimal plans. This sub-optimality arises because the optimizer’s decisions rely on cost models that are often inaccurate due to outdated statistics and simplifying assumptions (e.g., statistical independence of columns) (Stillger et al., 2001; Marcus et al., 2019; Chen et al., 2023). To exert more fine-grained control over plan selection, we propose embedding query hints (Marcus et al., 2021) within rewritten queries. In this work, we introduce a curated hint base and use LLMs to adaptively select and incorporate effective hints into rewritten queries.
@@ 6.2. Hint Injection
This section details our approach to using LLMs for identifying operations that could benefit from hints and then recommending appropriate ones. Unlike existing methods (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that recommend hints at the query level (based on global parameters like GUCs), where a single hint applies to all operations within that query, our strategy targets a more granular level. We found that query-level hints can be suboptimal, as a hint beneficial for one operation might negatively impact others. Therefore, we apply hints at a finer-grained level, such as to individual tables or specific operations within a query, to maximize positive impacts.
First, given the physical plan of $Q^{\prime}$, database semantics and data distribution statistics, we identify potential improvements that can be achieved through hints: (1) For every cardinality estimation in the plan, an LLM judges its reasonableness. If an estimation is deemed unreasonable, the LLM provides a recommended range and justification. (2) For all join operations, the LLM assesses whether the selected join operator (such as hash, merge, or nested-loop join) is suitable given the available statistics. Second, an LLM is then invoked to construct the most effective hints based on the obtained analysis and the query plan. It recommends join methods for specific tables and rectifies cardinality estimations. Regarding NO_MATERIALIZE, we apply this hint only if a CTE appears exactly once or processes a small, simple dataset. Finally, the proposed hints are then injected into $Q^{\prime}$, producing the final optimized query $Q^{*}$.
@@ 7.1. Experimental Set up
Testbed. Our experiments were conducted on a cloud server featuring a 32-core Intel Xeon Platinum 8352V CPU, 251 GB of RAM, a 942 GB SSD, and PostgreSQL v14.13. We utilize a NVIDIA A100-SXM4 GPU with 40 GB of memory to run the rewrite process of LLM-R2.
@@
  • LearnedRewrite (LR) (Zhou et al., 2021) uses a Monte Carlo Tree Search (MCTS) algorithm with a learned cost estimation model to explore a SQL policy tree of rewrite rule orders. We adopt the original implementation and use the cost estimation model provided in their official code repository.
  • LLM-R2 (Li et al., 2024) is a rule-based query rewrite system built on Apache Calcite (Begoli et al., 2018). It uses LLM’s In-Context Learning (ICL) ability to select query rewrite rules according to high quality demonstration queries.
@@ 7.2. Performance Comparison
The optimization benefits of QUITE over baselines are more pronounced on complex datasets. Notably, the improvement margin of QUITE over LR is substantially larger for the DSB dataset compared to the TPC-H and Calcite datasets. This disparity arises primarily because: (1) The LR we use is trained on the TPC-H, LR can proficiently identify effective rewrite rules and navigate their constrained search spaces. This existing proficiency limits the remaining optimization potential, thereby reducing the margin for QUITE to demonstrate substantial further improvements. (2) LR employs a greedy search within its MCTS algorithm, which proves less effective in the expansive search spaces of complex datasets like DSB, unlike its performance on simpler datasets (e.g., Calcite) with smaller search spaces and inherently less optimization potential.
@@ 7.3. Ablation Study
LLM-Agent-based Query Rewrite FSM To evaluate the impact of the FSM framework within our system, we performed two modifications: (1) the loop responsible for equivalence checking was removed, and (2) the decision stage was eliminated. These changes transformed the process into a linear, end-to-end pipeline while retaining all components and agent tools involved in query rewriting. At the same time, to further assess the impact of the MDP-based decision-making on the Reasoning agent’s capabilities, we conducted an ablation study by removing the MDP component. As shown in Table 5, QUITE outperforms its variant across all evaluation metrics. Notably, the average execution time with the MDP is reduced by 52.2% compared to the version without it, demonstrating that the MDP significantly enhances the effectiveness of the Reasoning agent.
@@
Query Hint Injection We evaluate the impact of hint injection on queries rewritten by QUITE. As shown in Table 3, the insertion of hints leads to a further decrease in average execution times by 1.8%, 3.8%, and 0.3% on TPC-H, DSB, and Calcite, respectively. These performance gains demonstrate that appropriate hints enable additional fine-grained optimization, even when queries have already been extensively optimized through rewriting. To highlight the advantage of our LLM-based hint injection approach, we compare its performance on rewritten queries against using LLMs to recommend Bao’s candidate hint sets. As shown in Table 7 and Figure 7, within the hint space defined at the GUC level provided by Bao, the overall trend indicates degraded performance.

### v2

@@ 4. LLM Agent-Based Query Rewriter
In this section, we present our LLM agent-based query rewriter and describe how to construct an efficient, robust query‐rewriting workflow using LLM agents.
@@ 4.1. MDP-based Reasoning Agent
Query rewrite is fundamentally a sequential decision problem (Finance and Gardarin, 1991b; Zhou et al., 2021), each transformation step changes the query structure and limits future optimization options. Reasoning LLM agents (e.g., DeepSeek-R1 (Guo et al., 2025)), capable of chain-of-thought planning and multi-step logical inference, appear well-suited to this task. However, our experiments show that they still produce 13.0% nonequivalent outputs and yield minimal execution improvements (Section 7.2). We observe that query rewrite mirrors a Markov Decision Process (MDP), where states are query forms, actions are rewrite steps, and rewards are cost reductions (Puterman, 1990; Liu et al., 2023; Wang, 2025). By framing rewrite in this way, we provide the reasoning LLM agent with a clear, stepwise optimization goal that guides each inference step. Therefore, we design an MDP-based reasoning agent to produce high-performing rewrites.
@@ 4.2. FSM-based Query Rewrite
Single-agent approaches often produce suboptimal or invalid rewrites because a single agent is responsible for managing the entire rewrite pipeline (e.g., propose rewrites, verify validity, and assess efficiency), which overloads the model’s context capacity and ultimately increases the risk of hallucination. Inspired by multi-agent frameworks that assign specialized agents to distinct tasks (Hong et al., 2023), we break down query rewrite into discrete stages, each is handled by a dedicated LLM agent. To ensure the multi-agent system’s stability and correctness (Han et al., 2024), we implement a Finite State Machine (FSM) to orchestrate stable stage transitions.
@@ 4.2.1. Query Rewrite Process Decomposition
In practice, we decompose the rewrite process into four stages: Reasoning, Verification, Decision and Termination, as illustrated in Figure 5. The Reasoning stage generates candidate SQL rewrites. The Verification stage validates syntax and semantic equivalence, correcting any errors identified. The Decision stage determines whether to continue or terminate the rewrite process. The Termination stage ends the rewrite process and outputs the rewritten SQL with rewrite proposals.
@@ 29       $t\leftarrow t+1$;
Rewrite Agent. After the Reasoning agent generates a chain, the Rewrite agent extracts proposals from each node, groups them into four categories matching our structured knowledge base, evaluates each proposal’s expected reward, and selects the top SQL candidate. Since this candidate may introduce new structural complexities, it cannot be used directly for verification. Therefore, the rewrite agent undertakes an SQL enhancement process to uncover additional rewrite opportunities. The prompt template for this process is shown below:
@@ 3. Apply other efficiency-preserving transformations
This template provides a concise, user-curated checklist of the effective SQL enhancement transformations. Guided by this checklist, the rewrite agent revisits the selected SQL candidate to identify any remaining optimization opportunities. If none are found, it returns the original query unchanged.
@@
  • Equivalence correction. Once syntax error is corrected, the agent submits the query pair $\langle Q_{\text{orig}},Q_{\text{new}}\rangle$ to verify the equivalence and correct nonequivalent queries.
@@
(2) Evaluation and Feedback Loop. Based on the generated report, the Decision agent gives a binary judge to decide whether the rewrite process should terminate or not. If True, the Decision agent terminates the FSM process; If False, it retrieves higher-confidence rewrite proposals from the knowledge base, then stores the current query, report, and retrieved knowledge in an agent shared buffer for the next FSM iteration. We deliberately avoid introducing knowledge base retrieval in the initial state, because reasoning models have input content limitations—excessive guidance information can reduce the generation freedom, potentially compromising the quality of reasoning proposal chain (Guo et al., 2025). Finally, the FSM enters the Termination stage and outputs the final rewritten query along with its comprehensive optimization report.
@@ 5.1. Structured Knowledge Base
Since DBAs encode numerous scenario-specific rewrite strategies that LLMs cannot learn from sparse examples and that fail to generalize across diverse query contexts, we consolidate these expert rules into a structured, scenario-indexed knowledge base. To this end, preparing the knowledge base involves the following steps: (i) collecting SQL rewrite knowledge from various resources, (ii) filtering untrustworthy content, (iii) enriching the entries, and (iv) retrieving the most relevant rules for each rewrite task. The Knowledge Base $\mathcal{K}$ is defined as follows:
@@
Step 3: Enhancing the knowledge. An LLM first summarizes questions and user’s responses from website Q&A units. These summaries are then augmented with corroborating information from official documents. To do this, we embed user response summaries using Sentence Transformers (Reimers and Gurevych, 2019). These embeddings are then used to query a pre-indexed knowledge base of extracted points, retrieving the top three matches via cosine similarity. Finally, an LLM evaluates these matches to confirm that they effectively explain the rewrite strategy underlying the Q&A unit. Suitable matches are incorporated, enriching Q&A units with concise explanations that aid future retrieval.
Step 4: Retrieving relevant knowledge. We classify these units into four categories: join optimization, constant folding, predicate simplification and others, which capture the essential aspects of query rewriting and provides a foundational structure for enabling precise retrieval of information. Since knowledge bases are often context-specific, we use the BM25 algorithm (Robertson et al., 2009) to retrieve the most relevant documents by computing the document relevance.
@@ 5.2. Hybrid SQL Corrector
Equivalence Error Correction. Verifying the equivalence between any pair of SQL queries is known to be an NP-hard challenge (Aho et al., 1979), and whether this problem admits a polynomial-time solution remains an open question. While numerous verification methods (Chu et al., 2017a; Wang et al., 2022, 2024; Ding et al., 2023) rooted in algebraic or symbolic reasoning have been proposed to prove equivalence, they typically operate on constrained query sets (e.g., specific operators, advanced SQL constructs). Recent studies have highlighted the potential of LLMs in deducing query equivalence with high confidence (Singh and Bedathur, 2024; Zhao et al., 2023). However, relying solely on LLMs carries inherent risks. Therefore, we integrate traditional verification tools and LLM capabilities in a two-stage process, forming a hybrid approach that achieves broader and more reliable equivalence judgments.
@@ 5.3. Agent Memory Buffer
Hash Join /*+ HashJoin(table table[ table...]) */ Forces the use of hash joins for specified tables.
@@
NOT MATERIALIZE /*+ NOT MATERIALIZE(table) */ Inlines the CTE to avoid materialization overhead.
@@ 6. Query Hint Recommender
Motivation. Query rewrite applies high-level transformations to give the DBMS optimizer a better starting point for plan generation. While this improved starting point is beneficial, the optimizer, which dictates the finer-grained execution operations, can still generate suboptimal plans. This sub-optimality arises because the optimizer’s decisions rely on cost models that are often inaccurate due to outdated statistics and simplifying assumptions (e.g., statistical independence of columns) (Stillger et al., 2001; Marcus et al., 2019; Chen et al., 2023). To exert more fine-grained control over plan selection, we propose embedding query hints (Marcus et al., 2021) within rewritten queries. In this work, we introduce a dedicated hint base and use LLMs to adaptively select and incorporate effective hints into rewritten queries.
@@ 6.2. Hint Injection
This section details our approach to using LLMs for identifying operations that could benefit from hints and then recommending appropriate ones. Unlike existing methods (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that recommend hints at the query level, where a single hint applies to all operations within that query, our strategy targets a more granular level. We found that query-level hints can be suboptimal, as a hint beneficial for one operation might negatively impact others. Therefore, we apply hints at a finer-grained level, such as to individual tables or specific operations within a query, to maximize positive impacts.
Building on the validated rewrite of $Q^{\prime}$, we now turn to refining its physical execution plan through targeted hint injection. First, given the physical plan of $Q^{\prime}$, database semantics and data distribution statistics, we identify potential improvements that can be achieved through hints: (1) For every cardinality estimation in the plan, an LLM judges its reasonableness. If an estimation is deemed unreasonable, the LLM provides a recommended range and justification. (2) For all join operations, the LLM assesses whether the selected join operator (such as hash, merge, or nested-loop join) is suitable given the available statistics. Second, an LLM is then invoked to construct the most effective hints based on the obtained analysis and the query plan. It recommends join methods for specific tables and rectifies cardinality estimations. Regarding NO_MATERIALIZE, we apply this hint only if a CTE appears exactly once or processes a small, simple dataset. Finally, the proposed hints are then injected into $Q^{\prime}$, producing the final optimized query $Q^{*}$.
@@ 7.1. Experimental Set up
Testbed. Our experiments were conducted on a server featuring a 32-core Intel Xeon Platinum 8352V CPU, 251 GB of RAM, a 942 GB SSD, and PostgreSQL v14.13. We utilize a NVIDIA A100-SXM4 GPU with 40 GB of memory to run the rewrite process of LLM-R2.
@@
  • LearnedRewrite (LR) (SIGMOD 2022)  (Zhou et al., 2021) uses a Monte Carlo Tree Search (MCTS) algorithm with a learned cost estimation model to explore a SQL policy tree of rewrite rule orders. We adopt the original implementation and use the cost estimation model provided in their official code repository.
  • LLM-R2 (VLDB 2025) (Li et al., 2024) is a rule-based query rewrite system built on Apache Calcite (Begoli et al., 2018). It uses LLM’s In-Context Learning (ICL) ability to select query rewrite rules according to high quality demonstration queries.
@@ 7.2. Performance Comparison
The optimization benefits of QUITE over baselines are more pronounced on complex datasets. Notably, the improvement margin of QUITE over LR is substantially larger for the DSB dataset compared to the TPC-H and Calcite datasets. This is primarily because: (1) The LR we use is trained on the TPC-H, LR can proficiently identify effective rewrite rules and navigate their constrained search spaces. This existing proficiency limits the remaining optimization potential, thereby reducing the margin for QUITE to demonstrate substantial further improvements. (2) LR employs a greedy search within its MCTS algorithm, which proves less effective in the expansive search spaces of complex datasets like DSB, unlike its performance on simpler datasets (e.g., Calcite) with smaller search spaces and inherently less optimization potential.
@@ 7.3. Ablation Study
LLM-Agent-based Query Rewrite FSM. To evaluate the impact of the FSM framework within our system, we performed two modifications: (1) the loop responsible for equivalence checking was removed, and (2) the decision stage was eliminated. These changes transformed the process into a linear, end-to-end pipeline while retaining all components and agent tools involved in query rewriting. At the same time, to further assess the impact of the MDP-based decision-making on the Reasoning agent’s capabilities, we conducted an ablation study by removing the MDP component. As shown in Table 5, QUITE outperforms its variant across all evaluation metrics. Notably, the average execution time with the MDP is reduced by 52.2% compared to the version without it, demonstrating that the MDP significantly enhances the effectiveness of the Reasoning agent.
@@
Query Hint Injection. We evaluate the impact of hint injection on queries rewritten by QUITE. As shown in Table 3, the insertion of hints leads to a further decrease in average execution times by 1.8%, 3.8%, and 0.3% on TPC-H, DSB, and Calcite, respectively. These performance gains demonstrate that appropriate hints enable additional fine-grained optimization, even when queries have already been extensively optimized through rewriting. To highlight the advantage of our LLM-based hint injection approach, we compare its performance on rewritten queries against using LLMs to recommend Bao’s candidate hint sets. As shown in Table 7 and Figure 7, within the hint space defined at the GUC level provided by Bao, the overall trend indicates degraded performance.